Let X and Y have joint probability mass function fX,Y (x, y) = (x + y)/30 for x = 0, 1, 2, 3 and y = 0,1,2. Find:

(a) Pr{X ≤ 2, Y = 1}(b) Pr{X > 2, Y ≤ 1}

(c) Pr{X +Y = 4}. (d) Pr{X > Y }.

(e) the marginal probability mass function of Y , and (f) E[XY].
Answer #1

(a)

PlXS 1 ) +1P ( X =10, 30-30 s 1 ) ( X Y 1 ) ( X + P = 1, + P 0+1= = 2, Y = 1)-. 30_ + 30

(b)

3+1 3+0 7 30 30 = _ = 0.2333

(c)

2+2 3+1 30 30 7 0.2333 30

(d)

18 30

(e)

Following table shows the joint and marginal pdfs:

X
0 1 2 3 P(Y=y)
0 0 1/30 2/30 3/30 6/30
Y 1 1/30 2/30 3/30 4/30 10/30
2 2/30 3/30 4/30 5/30 14/30
P(X=x) 3/30 6/30 9/30 12/30 1

The marginal pdf of Y is

Y P(Y=y)
0 6/30
1 10/30
2 14/30

(f)

Following table shows the calculations for E(XY):

X Y P(X=x,Y=y) xyP(X=x,Y=y)
0 0 0 0
0 1 1/30 0
0 2 2/30 0
1 0 1/30 0
1 1 2/30 2/30
1 2 3/30 6/30
2 0 2/30 0
2 1 3/30 6/30
2 2 4/30 16/30
3 0 3/30 0
3 1 4/30 12/30
3 2 5/30 30/30
Total 72/30

So,

72 E(XY) _ 〉 ·ryP(z, y) =-= 2.4 30